OpenStudy (anonymous):

The perimeter of a rectangle equals the sum of the lengths of its four sides. if the width of the rectangle is 8 inches and the perimeter is 52 inches. What is the length of the rectangle?

OpenStudy (anonymous):

8+8+x+x=52 where x=length

OpenStudy (anonymous):

2x=36 x=18

OpenStudy (anonymous):

8+8+x+x=2 16+2x=52 -16 -16 2x=36 /2 /2 x=18